b. You bought a sachet water machine from Indie Inc. The cost of the machine was GH¢35,000. At that time, you asked for the payment to be deferred, and a contract was written.You will then pay for the machine in a lump sum at the end of 2 years, with interest at a rate of 2% per quarter-year. According to the contract, if you ceased buying the packaging material from Indie Inc. at any time prior to 2 years, the full payment due at the end of 2 years would automatically become due. One year later, you decided to buy the packaging material elsewhere and stopped buying from Indie Inc., whereupon Indie Inc., per the contract terms, asked for the full payment that is due at the end of 2 years to be paid immediately. You were unhappy about this, so Indie Inc. offered as an alternative to accept the GH¢35,000 with interest at 10% per semiannual period for the 12 months that you had been buying the packaging material from Indie Inc. Which of the alternatives should you accept? Explain.
1
Expert's answer
2020-06-18T18:33:16-0400

The first option of seeing the contract run as per the terms agreed at the beginning.

This is because the alternative of paying the 35000 at an interest of 10% for the 12 months would cost more by 1,342.


Workings


"FV= PV(1+r)^{n}"


First option

"FV= 35000(1+0.02)^{8}"

"FV=41,008"


Second option

"FV=35000(1+0.1)^{2}"

"FV= 42,350"